# A New Dataset for Monocular Depth Estimation Under Viewpoint Shifts

### v1

@@ 6 Conclusion
In this paper, we addressed the underexplored problem of viewpoint shifts in monocular depth estimation for camera systems. We introduced a novel dataset collected using two dashcams in various positions and orientations. Our method leverages homography estimation and object detection as a cost-effective alternative to expensive lidar sensors for ground truth depth generation. We validated our approach on the KITTI dataset, demonstrating a high correlation (Spearman score of 0.97) with lidar measurements and measuring an absolute relative error comparable to the one using lidar. We showed how the intuitive limitations relative to the road geometry of the proposed approach are not impactful in practice on a public dataset like KITTI, and, by extension being the two geometries similar, to our dataset.
@@
Using the MonoViT model, we quantified the effects of viewpoint shifts on depth estimation performance. Our experiments revealed that certain camera positions significantly degrade accuracy. In particular combinations of pitch and roll and pitch and yaw have the most degrading effects. As a matter of fact, the most impactful shift induces an increased error of above 26 points in both zero-shot and fine tuning, which is more than double from the base position. Moreover, a study on inference scale of the model highlighted that camera pitch directly affects perceived scale and that, in general, performance degradation is correlated to scale distortion, suggesting that indeed automatic scaling of depth models at inference time could help mitigate the problem and thus be a promising research direction for generalization.
These results highlight the vulnerability of current depth estimation models to geometric shifts, particularly combinations of pitch and yaw angles. Fine tuning on the base camera data improved overall performance but did not fully mitigate the effects of viewpoint changes.
Our work provides insight into the limitations of current self-supervised monocular depth estimation models and emphasizes the importance of considering viewpoint variations in real-world applications. Future research should focus on developing more robust models that can generalize across different viewpoints, potentially by incorporating our dataset and evaluation methodology into training and testing pipelines. Additionally, expanding the dataset to include different vehicle types, camera models and more viewpoints would further enhance its utility.

### v3

@@ VI Conclusion
In this paper, we addressed the underexplored problem of viewpoint shifts in monocular depth estimation for camera systems. We introduced a novel dataset collected using two dashcams in various positions and orientations. Our method leverages homography estimation and object detection as a cost-effective alternative to expensive LIDAR sensors for ground truth depth generation. We validated our approach on the KITTI dataset, demonstrating a high correlation (Spearman score of 0.97) with LIDAR measurements and measuring an absolute relative error comparable to the one using LIDAR.
@@
Using the MonoViT model, we quantified the effects of viewpoint shifts on depth estimation performance. Our experiments revealed that certain camera positions significantly degrade accuracy. In particular, combinations of pitch and roll and pitch and yaw have the most degrading effects. As a matter of fact, the most impactful shift induces an increased error of above 26 points in both zero-shot and fine tuning, which is more than double the error of the base position. Moreover, a study on inference scale of the model highlighted that camera pitch directly affects perceived scale and that, in general, performance degradation is correlated to scale distortion, suggesting that indeed automatic scaling of depth models at inference time could help mitigate the problem and thus be a promising research direction for generalization.